# T-Explainer: A Model-Agnostic Explainability Framework Based on Gradients

### v1

@@ 6 Computational Performance
Another challenge that XAI methods have to face is related to computational performance. Shapley-based approaches like SHAP apply sampling approximations because the exact version of the method is an NP-hard problem with exponential computing time regarding the number of features [40]. SHAP is a leading method in feature attribution tasks due to its axiomatic properties, with a range of explainers based on different algorithms, from the exact version to optimized approaches. However, according to its documentation, the most precise algorithm is only feasible for modelings that are nearly limited to $15$ features [^2] https://shap.readthedocs.io/en/latest/generated/shap.ExactExplainer.html.
@@
Figure 3: Running times of SHAP and T-Explainer applied to the 3H-NN model trained on 16-dimensional synthetic data.
SHAP explainer was the fastest. However, the method was the most unstable, as we demonstrated in the stability assessments. The T-Explainer’s performance is close to KernelSHAP, while the SHAP exact explainer running time has grown exponentially. In such a performance test, we used a reduced dataset instead of the 20-FT data because applying the SHAP exact explainer in tasks with more than $16$ features is impossible. Most of the Machine Learning problems consider high-dimensional data. For example, the HIGGS is a 28-dimensional dataset with over $98$ thousand instances. With that dataset, it is impossible to apply the exact version of SHAP, and the optimized version of SHAP performs very poorly. Something similar could be said about the HELOC data, which recently motivated a challenge by FICO [26]. In this sense, T-Explainer is placed as a competitive XAI approach regarding computational resources.
@@ 7 Discussion, Limitations, and Future Work
We conducted experiments and comparisons using tabular data to create classification models for simplicity and because it is a typical setup for machine learning problems in academia and industry [15, 69]. We highlight that our experimental setup followed the assessment practices in recent literature [9, 33, 60, 11], focusing on Neural Networks as black boxes [38] and including datasets where leading XAI approaches have stability and applicability difficulties. However, we designed the T-Explainer considering more complex data representations, such as 2D images, videos, or semantic segmentation. We are extending our approach to encompass simplification masking approaches, a strategy widely used by many XAI methods [46, 40].
Another aspect under improvement is the $\mathbf{h}$ optimization module. As far as we know, there is no closed solution to determine an optimum value to $\mathbf{h}$ in this context. During our investigations, we observed that the finite difference computation presents a certain instability for instances near the model’s decision boundaries due to discontinuities imposed by the decision boundaries. The literature on numerical methods brings a number of alternatives to deal with discontinuities when approximating derivatives through finite differences [62, 51]. As part of the T-Explainer’s core, we are investigating alternatives to refine the $\mathbf{h}$ optimization process while maintaining computational efficiency.
A comprehensive assessment of a local explanation technique requires extensive quantitative [14] and qualitative [66] evaluations, including insights from human analysts [70]. We prioritized quantitative metrics in this work, particularly those concerning stability, but we acknowledge the value of humans in the loop assessment. Thus, we are committed to comprehensively validating the T-Explainer using metrics beyond those discussed here.
@@ 8 Conclusion
In this paper, we presented the T-Explainer, a Taylor expansion-based XAI technique. T-Explainer is a deterministic local feature attribution method built upon a solid mathematical foundation that grants it relevant properties such as local accuracy, missingness, and consistency. Moreover, T-Explainer is designed to be a model-agnostic method able to generate explanations for a wide range of machine learning models, as it relies only on the models’ outputs without accounting for their internal structure or partial results. The provided experiments and comparisons demonstrate that T-Explainer is quite stable and comparable to model-specific techniques, thus being a valuable explanation alternative.
@@ Acknowledgments
This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001, and in part by the São Paulo Research Foundation (FAPESP) – Finance Code #2022/09091-8. The views expressed are those of the authors and do not reflect the official policy or position of the sponsors.

### v2

@@ 6 Computational Performance
Another challenge that XAI methods have to face is related to computational performance. Shapley-based approaches like SHAP apply sampling approximations because the exact version of the method is an NP-hard problem with exponential computing time regarding the number of features [39]. SHAP is a leading method in feature attribution tasks due to its axiomatic properties, with a range of explainers based on different algorithms, from the exact version to optimized approaches. However, according to its documentation, the most precise algorithm is only feasible for modelings that are nearly limited to $15$ features [^2] https://shap.readthedocs.io/ExactExplainer, visited on July 2024.
@@
Figure 2: Running times of T-Explainer and SHAP explainers applied to the 3H-NN model trained on 16-dimensional synthetic data.
SHAP explainer was the fastest. However, the method was the most unstable, as we demonstrated in the stability evaluations. The T-Explainer’s performance is close to KernelSHAP, while the SHAP exact explainer running time has grown exponentially. In such a performance test, we used a reduced dataset instead of the 20-FT data because applying the SHAP exact explainer in tasks with more than $16$ features is impossible. Most of the Machine Learning problems consider high-dimensional data. For example, the HIGGS is a 28-dimensional dataset with over $98$ thousand instances. With that dataset, it is impossible to apply the exact version of SHAP, and the optimized version of SHAP performs very poorly. Something similar could be said about the HELOC data, which recently motivated a challenge by FICO [25]. In this sense, T-Explainer is placed as a competitive XAI approach regarding computational resources.
@@ 7 Discussion, Limitations, and Future Work
Another aspect under improvement is the $\mathbf{h}$ optimization module. As far as we know, there is no closed solution to determine an optimum value to $\mathbf{h}$ in this context. During our investigations, we observed that the finite difference computation presents a certain instability for instances near the model’s decision boundaries due to discontinuities imposed by the decision boundaries. The literature on numerical methods brings a number of alternatives to deal with discontinuities when approximating derivatives through finite differences [62, 50]. We refer to methods that develop Generalized Finite Differences (GFD) among the alternative approaches [57, 13]. As part of the T-Explainer’s core, we are investigating alternatives to refine the $\mathbf{h}$ optimization process while maintaining computational efficiency.
We conducted experiments and comparisons using tabular data to create classification models for simplicity and because it is a typical setup for machine learning problems in academia and industry [16]. We highlight that our experimental setup followed the evaluation practices in recent literature [10, 61, 12], focusing on Neural Networks as black boxes [37] and including datasets where leading XAI approaches have stability and applicability difficulties. However, we designed the T-Explainer considering more complex data representations, such as 2D images, videos, or semantic segmentation. We are extending our approach to encompass simplification masking approaches, a strategy widely used by many XAI methods [47, 39].
A comprehensive assessment of an explanation method requires quantitative [15] and qualitative [65] measures. We prioritized quantitative metrics in this work, particularly those for evaluating consistency, continuity, and correctness [42]. We are committed to comprehensively validating the T-Explainer using metrics beyond those discussed here.
@@ 8 Conclusion
In this paper, we presented the T-Explainer, a Taylor expansion-based XAI technique. It is a deterministic local feature attribution method built upon a solid mathematical foundation that grants it relevant properties such as local accuracy, missingness, and consistency. Moreover, T-Explainer is designed to be a model-agnostic method able to generate explanations for a wide range of machine learning models, as it relies only on the models’ outputs without accounting for their internal structure or partial results. The experiments and comparisons demonstrate that T-Explainer is stable and accurate compared to model-specific techniques, thus making it a valuable explanation alternative.
@@ Acknowledgments
This work was supported in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior–Brasil (CAPES)–Finance Code 001, and in part by São Paulo Research Foundation (FAPESP) under Grant 2022/09091-8. The opinions, hypotheses, and conclusions or recommendations expressed in this material are the authors’ responsibility and do not necessarily reflect the views of the funding agencies.